# Inducing Group Fairness in LLM-Based Decisions

### v1

@@ E.1 Prompting
E.2 In-Processing
@@
Min Diff Prost et al. (2019) is an approach that has been successful in remediating to achieve equality of opportunity in the ‘classical’ setting. The central insight behind the Min Diff approach is that the distance between the probability of a false positive for instances from groups and majority can be included in the loss. This encourages those distributions to be closer together, which in turns pushes the false positive rates for group and majority towards each other.
In the equations that follow, let $(x,y)$ represent a feature and a label, where $x\in\mathcal{X}$, and $y\in\{0,1\}$.
We use the following loss:
$L_{\textit{IP}}=L_{\textit{CE}}(\hat{Y},Y)+\lambda\hskip 2.168pt\textit{MMD}\left(\hat{Y}|Y=0,G=0;\hat{Y}|Y=0,G=1\right).$ (2)
Where $L_{\textit{IP}}$ is the in-processing loss, $L_{\textit{CE}}$ is the usual cross entropy loss, MMD is a Maximum Mean Discrepancy kernel that gives the distance between the distributions of probability of a false positive for group and associated majority, and $\lambda$ is a parameter that trades off between the two loss terms (and thus between performance and fairness).
@@
In order to adapt this to the LLM case, we use the Min Diff loss during fine tuning. This approach does not work in the zero-shot case where no fine tuning is performed, but the following approach does.
E.3 Post-Processing
@@
Recent work has demonstrated how in-processing techniques can be adapted to post-processing scenarios Ţifrea et al. (2024). We leverage this approach to fit a post-processed ‘emfairening’ model. The emfairening model’s predictions are added to the unremediated models predictions in logit space:
$P_{pp}(y|x)=\sigma(z(P_{bl}(y|x))+z(P_{ef}(y|x))),$ (3)
@@
where $\sigma$ is the logistic function, $z$ is the logit function, $P_{bl}(y|x)$ is the baseline prediction distribution, $P_{ef}(y|x)$ is the emfairening model’s distribution, and $P_{pp}(y|x)$ gives the combined post-processed distribution. The emfairening model can be trained with the following loss:
[…]
@@
where KL is the Kullback-Leibler divergence and MMD is the Maximum Mean Discrepancy kernel. As with the Min Diff approach, the Maximum Mean Discrepancy kernel pushes the probabilities of a false positive for group and majority (and thus false positives rates for group and majority) toward each other. The KL divergence term prevents ‘catastrophic forgetting’ in the emfairening model; that is, we encourage the emfairening model to not stray too far from the performant baseline model. This ensures that we maintain acceptable classifier performance when making emfairened predictions. As with the in-processing method, $\lambda$ trades off between the two loss terms and thus dials between fairness and performance.
@@
Note that we are free to choose the baseline model $P_{bl}(y|x)$. As such, we can apply this approach directly to the LLM in a zero-shot environment or apply it to a fine-tuned model.
Appendix F Full Benchmark Ratio Gap Results
Group Zero-Shot FPR Ratio Fine-Tuned FPR Ratio
@@ Appendix F Full Benchmark Ratio Gap Results
Table 2: False positive rate ratios that quantify the magnitude of violation of equality of opportunity for a classifier built on PaLM 2.
We only report the ratio gaps for the two groups with the highest gaps (Jewish and Muslim) in Tab. 1 in Sec. 4. The full table is given in Table 2
Appendix G Use of Scientific Artifacts

### v2

@@ 5.1 Prompting
5.2 In-Processing
Let $(x,y)$ represent an input prompt and a label, where $x\in\mathcal{X}$, and $y\in\{0,1\}$. We use the following loss:
$L_{\textit{IP}}=$ $L_{\textit{CE}}(\hat{Y},Y)+\lambda D\left(\hat{Y};G|Y\!=\!0\right),$ (2)
Where $L_{\textit{IP}}$ is the in-processing loss, $L_{\textit{CE}}$ is the usual cross entropy loss for learning a classification task, and $D$ is a statistical divergence promoting the decision $\widehat{Y}$ to be independent of the sensitive attribute $G$.
@@ 5.2 In-Processing
Among many choices for the statistical divergence, we focus on Min Diff Prost et al. (2019), which is an approach that has been successful in remediating to achieve equality of opportunity in the ‘classical’ setting in industrial settings Prost and Beutel (2020). The central insight behind the Min Diff approach is that the distance between the probability of a false positive for instances from groups and majority can be included as the loss. This encourages those distributions to be closer together, which in turns pushes the false positive rates for group and majority towards each other. In particular, Min Diff uses MMD as a Maximum Mean Discrepancy kernel that gives the distance between the distributions of probability of a false positive for group and associated majority, and $\lambda$ is a parameter that trades off between the two loss terms (and thus between performance and fairness).
@@
To adapt Min Diff to the LM decisions, we use the Min Diff loss during fine tuning. This approach cannot work in the zero-shot case where no fine tuning is performed.
5.3 Post-Processing
@@
Recent work has demonstrated how in-processing techniques can be adapted to post-processing scenarios Ţifrea et al. (2024). We leverage this approach to fit a post-processed ‘emfairening’ model. The emfairening model’s predictions are added to the unremediated models predictions in logit space, i.e., for all prompts $x$ and label $y$:
$\pi_{pp}(y|x)\propto\pi_{\text{ref}}(y|x)\pi_{\text{emf}}(y|x),$ (3)
@@
where $\pi_{\text{ref}}(y|x)$ is the baseline prediction distribution, $\pi_{\text{emf}}(y|x)$ is the emfairening model’s distribution, and $\pi_{pp}(y|x)$ gives the combined post-processed distribution. The emfairening model can be trained with the following loss:
$L_{PP}=$ $\textit{KL}\left(P_{pp}\|P_{ref}(y|x)\right)+\lambda D\left(\hat{Y};G|Y\!=\!0\right)$ (4)
@@
where KL is the Kullback-Leibler divergence and $D$ is a fairness promoting divergence. In this paper, We use the industry trusted MMD kernel Prost and Beutel (2020). The KL divergence term prevents ‘catastrophic forgetting’ in the emfairening model; that is, we encourage the emfairening model to not stray too far from the performant baseline model. This ensures that we maintain acceptable classifier performance when making emfairened predictions. As with the in-processing method, $\lambda$ trades off between the two loss terms and thus dials between fairness and performance.
We note that the post-processing formulation here resembles controlled decoding methods used to steer the generation of language models towards high reward outcomes Mudgal et al. (2024) with a reward that captures the fairness of the outcome.
@@
We also note that we are free to choose the baseline model $\pi_{\text{ref}}(y|x)$. As such, we can apply this approach directly to any LM in a zero-shot manner or apply it to a fine-tuned model.
6 Experiment Results
Group fairness without remediation.
First, we evaluate the two classifier approaches on PaLM 2 S Anil et al. (2023) with respect to equality of opportunity. These experiments follow the methodology described in Sec. 3 and use the Civil Comments Identity Borkan et al. (2019) dataset.
Group Prompt-based Classifier Embedding-based Classifier
Muslim 1.89 2.24
Jewish 1.48 1.71
Table 1: False positive rate ratios that quantify the magnitude of EO violation for a classifier built on PaLM 2 S. We only include the two groups with the highest gaps. Here we used ‘Christian’ as the majority group.
The FPR ratios for the two groups with the highest ratio gaps are given in Tab. 1 and the results for all groups are given in Tab. 2 (Appendix). These elevated FPR ratios imply a need for group fairness remediation. In the remainder of this section, we will focus on remediation approaches and analyze their empirical performance vs fairness trade-offs.
Remediation of prompt-based classifiers
Muslim Group     Jewish Group
(a) Prompt-based Classifiers
(b) Embedding-based Classifiers
(c) Model Transfer
Figure 2: Pareto frontiers of different remediation techniques. The left plot shows the performance and fairness of prompt-based classifiers, and the middle plot of embedding-based classifiers. The unremediated classifier setting is denoted by a ‘+’ and prompting-based remediation methods are denoted by single symbols. Note that the in-processing baseline is inapplicable to prompt-based classifiers. Each point for in-processing and post-processing is generated by setting different values for $\lambda$ in Equations (2) and (4) in the appendix. The dashed and solid lines give the Pareto frontier where performance can only be gained by sacrificing fairness, for post-processing and in-processing, respectively. The right plot gives the effect of model transfer. We fit a post-processing remediation model to the PaLM 2 S model then compare the effects of applying it to PaLM 2 S (native) versus the larger PaLM 2 L model (transfer). The lines give the Pareto frontier (solid for native and dashed for transfer).
We start with analyzing remediation techniques in prompt-based classifiers. Results are shown in Fig. 2(a). For the post-processing method, each point in the plot is generated by varying the regularizer strength (see Appendix 5.3 for more details).
We observe that the post-processing method improve fairness without severely degrading the performance of the classifier. In contrast, prompt-based remediations either increase performance while keeping the high FPR ratio, or reduce the FPR ratio at the cost of performance.
Remediation of embedding-based classifiers.
We compare the Pareto frontiers of fairness and performance for each remediation technique in Fig. 2(b). As before, each point of regularization-based techniques in the plot is generated by varying the regularizer strength that trades off between performance and fairness terms in the objective function.
There are a few takeaways from these experiments. First, embedding-based classifiers show superior performance compared to prompt-based classifiers (see Fig. 2(a)). Second, as before, we observe improved group fairness for regularization-based methods without significantly degrading the performance of the classifier. However, the in-processing technique generally performs better than the post-processing technique in this setting. Finally, prompt-based remediation methods show some fairness and performance benefit but are generally less controllable and less effective than in-processing and post-processing methods.
Transfer of remediation to an unseen model.
In this experiment, our goal is to apply the remediation to a grey-box model which only provides access to logits (and not the last layer activations), i.e., a prompted LM. As such, we use the Google News 128-dimensional embedding model Bengio et al. (2000) [^1] https://www.kaggle.com/models/google/nnlm. to embed the query and use these embeddings for any subsequent remediation. Our setup enables us to operate in environments where drawing embeddings from the LM is not feasible. One interesting case is where we train a post-processing remediation model on one LM, and then apply it to remediate another LM (with grey-box access); can we reuse the existing fairness model to improve fairness with the new model?
Fig. 2(c) gives the results this model transfer learning scenario. Note also that the Pareto frontier achieved by the smaller PaLM 2 S model in Fig. 2(c), where Google News embeddings are used, is just slightly degraded from the Pareto frontier given in the upper plot of Fig. 2(a) where model activiations are used as embeddings, making this a promising approach for fairness remediation. Importantly, we find that post-processing model is still able to improve fairness when transferred (although this comes at the cost of higher performance degradation than when applied to the model that it was trained on). This suggests that we may be able to train universal fairness mitigation heads that could be applied to any LM with grey-box access and provide fairness benefits for classification tasks.
7 Concluding Remarks
Fairness is an important consideration for classifiers in industry. Even though the research community has made significant progress on training fair classifiers, the recent shift towards prompt-based classifiers requires exploring new fairness solutions for prompt-based decision making.
We study the group fairness of two classes of LM-based classifiers. We identify that LM-based classifiers may exhibit group unfairness. We introduce and evaluated three remediation techniques to improve fairness while maintaining acceptable performance for LM-based classifiers. We find that prompt-based techniques offer limited benefit and are in general outperformed by in-processing and post-processing techniques.
Within the scope of the evaluated prompts, we find that embedding-based classifiers exhibit superior performance compared to “out-of-the-box” prompt-based classifiers in our setup. In addition, we find that the in-processing method consistently provides favorable performance vs fairness trade-off on embedding-based classifiers. We conclude that for remediating an embedding-based classifier, in-processing is a more robust approach. In other LM-based classification settings where in-processing cannot be applied (prompt-based classifiers and transfer tasks) the post-processing technique provides promising results.
We find that the prompt-based remediation methods have little to no impact of prompts on fairness, while counter-intuitively, we observe that fairness-oriented prompts may slightly improve performance in some cases for the less specific ‘Please be Fair’ (PBF) and ‘Please be Fair to Super Group’ (PBF2SG) methods. This is not surprising given that fairness is a distributional issue, and hence prompting may not necessarily provide the distribution matching effects that we expect from remediation.
Appendix A Limitations
  • We find that prompting-based remediation methods are less flexible and effective than in-processing and post-processing methods. However, we do not make an exhaustive search of possible prompts and other researchers may find prompting-based remediation methods that work. Furthermore, it has been observed that the capabilities of language models improve with the model size Wei et al. (2022b), and this could have a beneficial effect on prompt-based method effectiveness as LMs become larger and more capable.
  • Apart from prompting-based remediation methods, the proposed remediation techniques require grey-box access to the logits of the model prior to sampling, and may not be applicable if the model only provides black-box access.
  • Our experiments are focused on equal opportunity (EO) notion of group fairness. There is no guarantee that they will generalize to other notions of fairness, and, importantly, their application does not imply that a classifier is abstractly fair, as all different notions of group fairness have their own limitations and might even be at odds with each other.
  • LM-based classifier inference is very expensive when compared to simpler models, and the performance of LM-based classifiers does not yet justify that cost (for example, our LM-based classifiers are less effective than baseline methods given by the authors of the Civil Comments dataset paper Borkan et al. (2019)). An implicit assumption of this work is that the performance of LM-based classifiers will improve enough over time to justify their high inference costs and become deployed systems where fairness considerations are in play.
  • We considered only one language model (PaLM 2) and one dataset (Civil Comments Identity) in English. So it remains to be seen how much our findings generalize. Having said that, given that we already find performance disparities across subgroups in this limited case, the need for developing fairness remediation techniques for LM-based decision making systems is justifiably real.
  • We only experiment with a few-handcrafted prompts for classification, and did not compare against chain-of-thought Wei et al. (2022c), self-consistency Wang et al. (2022), and automated prompt generation Gao et al. (2021) techniques as adapting them to induce group fairness was not trivial and is left for future work.
  • We do not benchmark other popular techniques, such as low-rank adaptation Hu et al. (2021), prompt-tuning Lester et al. (2021), and other parameter-efficient fine-tuning techniques Liu et al. (2022) for the in-processing method.
Appendix B Risks
There are three risks we would like to call out:
Appendix C Data and Data Processing
The dataset contains 405,130 training instances, 21,293 validation instances, and 21,577 test instances. We make use of all three splits in our work. The training set was used for training, validation set for classifier threshold selection, and the test set for reported results.
Appendix D Full Benchmark Ratio Gap Results
Group Prompt-based Classifier (FPR Ratio) Embedding-based Classifier (FPR Ratio)
@@ Appendix D Full Benchmark Ratio Gap Results
Table 2: False positive rate ratios that quantify the magnitude of violation of equality of opportunity for an unremediated classifier built on PaLM 2 (Anil et al., 2023)
We only report the ratio gaps for the two groups with the highest gaps (Jewish and Muslim) in Tab. 1 in Sec. 6 for the unremediated case. The full table is given in Table 2
Appendix E Use of Scientific Artifacts